OpenStudy (anonymous):

Could someone explain me the following integral step by step using trigonometric sustitution? Indefinite integral

OpenStudy (anonymous):

\[\int\limits_{a}^{b}\sqrt{9-x^{2}}\]

OpenStudy (anonymous):

I tried it by myself and wolfram and I are different in just 1 number, he's dividing a part by 2 and I'm not, so help would be appreciated, I can take a picture of my procedure.

OpenStudy (chaise):

Recall that when you integrate you divide by the derivative of the inner function, in this case, the derivative of the inner function is x^2, whose derivative is 2x. Wolfram alpha might also have an extra x in there too, that you are missing.

OpenStudy (anonymous):

The deal is that I was asked to do this using trigonometric sustitution for my homework, and since this is the first time I integrate using this method I need some explanation heh.

OpenStudy (chaise):

Well, you need to make this substution: \[\sqrt{a^2-x^2} --> asin(x)\] In this case, your a will be 3.

OpenStudy (chaise):

Try watching this video :) https://www.youtube.com/watch?v=3lC5AuCFK4c

OpenStudy (anonymous):

I got it right already, thank you!